joe can paint the barn in 11 hours by himself. Steve has never timed himself painting a barn before so we don't know how fast he can complete the job. Working together, the two finish the job in 7 hours. How long would it take steve to paint the barn alone?
Answer by TimothyLamb(4379)

x = hours Steve takes to paint the barn alone
---
rate:
r = w/t
---
rates are additive:
1/11 + 1/x = 1/7
---
1/x = 1/7 - 1/11
x = 1/(1/7 - 1/11)
1/(1/7 - 1/11)
---
copy and paste the above expression in to this calculator:
http://sooeet.com/math/scientific-calculator.php
---
x = 19.25
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---